2.5 44) Geometry.

The lenght of a recetangular playing field is 5 feet less than twice its width. If the perimeter of the playing dield is 230 ft, find the lenght and width of the field.

Answer by checkley71(8403)

LENGTH IS =2W-5 WIDTH IS =W THE THE PERIMETER IS 2(2W-5)+2W=230 THUS
4W-10+2W=230 6W=240 OE W=240/6 OR W=40 THEN THE LENGTH IS 2*40-5 OR 80-5 OR 75
PROOF 2*75+2*40=150+80=230